OpenStudy (anonymous):

(x-5)(2x+7)=0 solve equation

OpenStudy (anonymous):

Simplifying (x + -5)(2x + 7) = 0 Reorder the terms: (-5 + x)(2x + 7) = 0 Reorder the terms: (-5 + x)(7 + 2x) = 0 Multiply (-5 + x) * (7 + 2x) (-5(7 + 2x) + x(7 + 2x)) = 0 ((7 * -5 + 2x * -5) + x(7 + 2x)) = 0 ((-35 + -10x) + x(7 + 2x)) = 0 (-35 + -10x + (7 * x + 2x * x)) = 0 (-35 + -10x + (7x + 2x2)) = 0 Combine like terms: -10x + 7x = -3x (-35 + -3x + 2x2) = 0 Solving -35 + -3x + 2x2 = 0 Solving for variable 'x'. Factor a trinomial. (-7 + -2x)(5 + -1x) = 0 Subproblem 1 Set the factor '(-7 + -2x)' equal to zero and attempt to solve: Simplifying -7 + -2x = 0 Solving -7 + -2x = 0 Move all terms containing x to the left, all other terms to the right. Add '7' to each side of the equation. -7 + 7 + -2x = 0 + 7 Combine like terms: -7 + 7 = 0 0 + -2x = 0 + 7 -2x = 0 + 7 Combine like terms: 0 + 7 = 7 -2x = 7 Divide each side by '-2'. x = -3.5 Simplifying x = -3.5 Subproblem 2 Set the factor '(5 + -1x)' equal to zero and attempt to solve: Simplifying 5 + -1x = 0 Solving 5 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '-5' to each side of the equation. 5 + -5 + -1x = 0 + -5 Combine like terms: 5 + -5 = 0 0 + -1x = 0 + -5 -1x = 0 + -5 Combine like terms: 0 + -5 = -5 -1x = -5 Divide each side by '-1'. x = 5 Simplifying x = 5 Solution x = {-3.5, 5}

OpenStudy (nathan917):

*was just going to copy and past that :D* *smacked myself*

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

two solutions for this one

OpenStudy (anonymous):

x = {-3.5, 5}

OpenStudy (anonymous):

rofl... (x-5) = 0 or (2x-7) = 0 x =5 or x = 3.5

OpenStudy (nathan917):

x = {-3.5, 5} Why read his.! *He beat me again to it.* *smacked myself again* *harder this time*

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):